# Identifying Equivalent Training Dynamics

### v2

@@ Abstract
Study of the nonlinear evolution deep neural network (DNN) parameters undergo during training has uncovered regimes of distinct dynamical behavior. While a detailed understanding of these phenomena has the potential to advance improvements in training efficiency and robustness, the lack of methods for identifying when DNN models have equivalent dynamics limits the insight that can be gained from prior work. Topological conjugacy, a notion from dynamical systems theory, provides a precise definition of dynamical equivalence, offering a possible route to address this need. However, topological conjugacies have historically been challenging to compute. By leveraging advances in Koopman operator theory, we develop a framework for identifying conjugate and non-conjugate training dynamics. To validate our approach, we demonstrate that it can correctly identify a known equivalence between online mirror descent and online gradient descent. We then utilize it to: identify non-conjugate training dynamics between shallow and wide fully connected neural networks; characterize the early phase of training dynamics in convolutional neural networks; uncover non-conjugate training dynamics in Transformers that do and do not undergo grokking. Our results, across a range of DNN architectures, illustrate the flexibility of our framework and highlight its potential for shedding new light on training dynamics.
@@ 1 Introduction
The analysis and experimentation of deep neural network (DNN) training continues to uncover new – and in some cases, surprising – phenomena. By changing the architecture, optimization hyper-parameters, and/or initialization, it is possible to identify regimes in which DNN parameters evolve along trajectories (in parameter space) with linear dynamics [1, 2], low-dimensional dynamics [3], correlated dynamics [4], lazy and rich dynamics [5, 6], and oscillatory dynamics [7, 8]. In some cases, the training dynamics have been linked with the performance of the trained model [7, 9, 10], providing new insight in DNN generalization. Additionally, detailed understanding of the dynamics has led to improvements in training efficiency [4, 11], demonstrating the practical implications such work can provide.
To obtain a complete picture of DNN training, it is necessary to have a method by which equivalent dynamics can be identified and distinguished from other, non-equivalent dynamics. The construction of equivalence classes, which has fundamentally shaped the study of complex systems in other domains (e.g., phase transitions [12], bifurcations [13], defects in materials [14]), would advance the understanding of how architecture, optimization hyper-parameters, and initialization shape DNN training and could be leveraged to search for new phenomena. However, identifying equivalent and non-equivalent training dynamics is challenged by the need for methods that:
  • Go beyond the coarseness of loss. While useful as metrics, training and test loss can be shaped by non-dynamical features of the training (e.g., different initializations, different number of hidden units). Thus different losses are neither necessary nor sufficient to conclude non-equivalent training dynamics.
  • Respect permutation symmetry. DNNs are invariant to the re-ordering of their hidden units [15]. Thus, identifying that DNN parameters evolve along trajectories that occupy distinct parts of parameter space is not sufficient to conclude non-equivalent dynamics [16, 17].
We propose to use topological conjugacy [18], a notion of dynamical equivalence developed in the field of dynamical systems theory (Sec. 3.1), to address these limitations. Historically, topological conjugacy has been difficult to compute [19], especially when the equations governing the dynamical systems under study are not know. However, recent advances in Koopman operator theory [20, 21, 22] (Sec. 3.2) have enabled the identification of topological conjugacies from data [23] (Sec. 3.3). We explore the potential of this Koopman-based approach for identifying topological conjugacies in the domain of DNN training, finding that it is able to:
  • Recover a known nonlinear topological conjugacy, between the training dynamics of online mirror descent and online gradient descent [24, 25, 26] (Sec. 4.1);
@@
  • Uncover non-conjugate training dynamics between Transformers that do and do not undergo delayed generalization (i.e., “grokking”) [28, 29] (Sec. 4.4).
That the same framework can be used across a number of DNN architectures to study a variety of dynamical phenomena during training demonstrates the generality of this approach. We conclude by discussing how it can be further improved to enable greater resolution of equivalent dynamics, and how it can be used to shed greater light on DNN training (Sec. 5).
@@ 2.1 Identification of DNN training dynamics phenomena
Analytical results have been obtained for the DNN training dynamics of shallow student-teacher [30, 31] and infinite width [1, 2] networks. For modern architectures (e.g. CNNs, Transformers), the training dynamics have been probed via analysis of computational experiments. Application of dimensionality reduction has led to the observation that parameters are quickly constrained to being optimized along low-dimensional subspaces of the high-dimensional parameter space [3, 4]. Inspection of losses, parameter and gradient magnitudes, etc. led to the identification of several transitions in the training dynamics of CNNs during the initial several epochs [27]. While insightful, this prior work cannot – except at a coarse-grained level – be used to determine whether the dynamics associated with training different DNN models (or training the same DNN model with different choices in hyper-parameters or initialization) are equivalent.
@@ 2.2 Koopman operator theory applied to DNN training
Data-driven implementations of Koopman operator theory have been used to model the dynamics of DNN training [32, 33, 34]. Because of the linearity of the learned Koopman models (Sec. 3.2), using them in place of standard gradient-based methods has led to reductions in computational costs associated with training. Koopman-based methods have additionally been used to meta-learn optimizers for DNNs [35, 36]. The ability of Koopman models to capture features of training dynamics has been leveraged to develop new methods for pruning DNN parameters [37, 38] and new adaptive training methods [39]. While this prior work has demonstrated that accurate Koopman operator representations of the nonlinear training dynamics can be extracted, none have yet to utilize the theory to identify topological conjugacies.
@@ 3.1 Topological conjugacy
Given two discrete-time dynamical maps [^1] For sake of space, we describe only discrete-time dynamical systems, but the theory extends to continuous time dynamical systems. $T_{1}:X\rightarrow X$ and $T_{2}:Y\rightarrow Y$, a natural question to ask is whether they induce equivalent dynamical behavior. There are various possibilities for defining equivalence, but dynamical systems theory has made use of the notion of topological conjugacy [18] to identify when a smooth invertible mapping can be used to transform trajectories of $T_{1}$ to those of $T_{2}$ (and vice versa). Formally, $T_{1}$ and $T_{2}$ are said to be topologically conjugate if there exists a homeomorphism, $h$, such that
@@
It is straightforward to identify and construct conjugacies for linear systems. Let $X=Y=\mathbb{R}^{n}$, and let $T_{1}=A$ and $T_{2}=B$, where $A,B\in\mathbb{R}^{n\times n}$. These describe linear dynamical systems, as $x(t+1)=Ax(t)$ and $y(t+1)=By(t)$. In this setting, $A$ and $B$ are conjugate if there exists an $H\in\mathbb{R}^{n\times n}$, such that $y(t)=Hx(t)$ and $A=H^{-1}BH$. This can happen if and only if the eigenvalues of $A$ are the same as the eigenvalues of $B$. Thus, for linear systems, topological conjugacy can be used to construct equivalence classes, partitioning the space of dynamical systems into families of matrices that have the same spectra. However, for nonlinear systems, it is challenging to prove the existence or non-existence of conjugacies, limiting the use of topological conjugacy as a tool. In addition, historically it has not been possible to compute topological conjugacies for systems where the underlying dynamics are unknown.
@@ 3.2 Koopman mode decomposition
Figure 1: Schematic of Koopman operator theory based identification of conjugate dynamical systems. (A) By lifting nonlinear dynamics from a finite dimensional state-space to an infinite dimensional function space, a linear representation can be achieved (from which a finite dimensional approximation can be obtained). (B) The linearity of the Koopman operator enables a mode decomposition, which includes Koopman eigenvalues (orange), eigenfunctions (green), and modes (blue). (C) Dynamical systems with the same Koopman eigenvalues are topologically conjugate.
@@
An important feature of the Koopman eigenvalues is that they are invariant to permutations of the ordering of state-space variables. Consider $x=[x_{1},...,x_{n}]$ and $\tilde{x}=[x_{\sigma(1)},...,x_{\sigma(n)}]$, where
$\sigma:\{1,...,n\}\rightarrow\{1,...,n\}$ is a permutation. That is, $\tilde{x}$ is equivalent to $x$ via a re-ordering of its labels. In this case, the action of the Koopman operator is given by
$U^{t}g(\tilde{x})=\sum_{i=1}^{N}\lambda_{i}^{t}\phi_{i}(\tilde{x})v^{\prime}_{i}.$ (5)
Thus, the Koopman eigenvalues (and eigenfunctions) are the same as they were for the non-permuted system.
Many numerical methods for approximating the KMD from data have been developed and this has enabled its application as a tool for spatio-temporal decomposition that has been successfully applied to provide insight into complex, real-world dynamical systems [41, 42, 43, 44, 45]. Dynamic mode decomposition (DMD) [46, 47], the most popular of these methods, has spawned many variants [48, 49, 50, 51, 52]. In general, DMD-based approaches collect $T+1$ snapshots of data $x\in\mathbb{R}^{n}$, construct data matrices $Z=[x(0),...,x(T-1)]$ and $Z^{\prime}=[x(1),...,x(T)]$, where $Z,Z^{\prime}\in\mathbb{R}^{n\times(T+1)}$, and then approximate the Koopman operator by
@@
where $\dagger$ denotes the pseudo-inverse. Utilizing dictionaries with nonlinear functions [48] and time-delays [42, 43, 49] has led to improved results, further demonstrating how usage of the underlying Koopman operator theory can enhance the capture of complex dynamics.
@@ 3.3 Equivalent Koopman spectra implies topological conjugacy
Given that KMD provides a linear representation of nonlinear dynamical systems, identifying topological conjugacies through matching eigenvalues again becomes viable. Indeed, it has been proven that two discrete-time dynamical maps $T_{1}:X\rightarrow X$ and $T_{2}:Y\rightarrow Y$, each in the basin of attraction of a stable fixed point, are topologically conjugate if and only if the Koopman eigenvalues of the associated Koopman operators, $U_{1}$ and $U_{2}$, are the same [23] (Fig. 1C). That is, a topological conjugacy exists if and only if
@@
where $\lambda^{(1)}$ and $\lambda^{(2)}$ correspond to the eigenvalues associated with $U_{1}$ and $U_{2}$, respectively, and $N$ is the number of Koopman modes.
Computing the KMD from data is unlikely to yield the same exact Koopman eigenvalues for conjugate dynamical systems, due to the presence of noise and finite sampling. Therefore, a method for computing the distance between eigenvalues is necessary when making comparisons. Here, we make use of the Wasserstein distance [53, 54], a metric developed in the context of optimal transport that quantifies how far apart two distributions are. In the case where a small, finite number of Koopman eigenvalues are computed (which can be achieved, even for systems with a large number of observables, by performing dimensionality reduction or residual based pruning of modes [51]), the Wasserstein distance can be efficiently computed by using linear sum assignment.
@@ 4.1 Identifying conjugate optimizers
Figure 2: Conjugacy between online mirror descent and online gradient descent is identifiable from Koopman spectra. (A) Comparing example trajectories of variables optimized via OMD ($x_{1},x_{2}$) and OGD ($u_{1},u_{2}$), the existence of a conjugacy is not obvious. (B) Similarly, the existence of a conjugacy is not apparent when looking at the loss incurred by using OMD and OGD. (C) Comparing the Koopman eigenvalues associated with optimizing using OMD and OGD correctly identifies the existence of a conjugacy. The function optimized is in all subfigures is $f(x)=x^{4}$.
We compute the KMD associated with optimization using OMD and OGD by considering trajectories from many initial conditions, and compare the resulting Koopman eigenvalues (Appendix A.3). We find high overlap between the two spectra (Fig. 2C). Additionally, the two sets of eigenvalues have the same structure,. Namely, they consist only of real, positive eigenvalues. Similar results are found when applying KMD to OMD and OGD optimization of a different function (Fig. S1). Collectively, these results demonstrate that the Koopman-based spectral identification of topological conjugacies can successfully recover a known equivalence and provide support that it can be used to uncover equivalences in DNN training dynamics.
4.2 Identifying effect of width on fully connected neural network training
To start exploring the potential of our framework for identifying topological conjugacies in DNN training, we begin with a small-scale example. Namely, we consider a fully connected neural network (FCN) with only a single hidden layer, trained on MNIST (Appendix B.1). Consistent with other architectures, we find that the wider the FCN (i.e., the more units in the hidden layer), the better the performance and the lower the loss (Fig. 3A). Whether this is due to an increase in capacity, with more hidden units enabling a more refined solution, or whether this is due to a change in the training dynamics, leading to a better traversal of the loss landscape, is unclear.

### v3

@@ Abstract
Study of the nonlinear evolution deep neural network (DNN) parameters undergo during training has uncovered regimes of distinct dynamical behavior. While a detailed understanding of these phenomena has the potential to advance improvements in training efficiency and robustness, the lack of methods for identifying when DNN models have equivalent dynamics limits the insight that can be gained from prior work. Topological conjugacy, a notion from dynamical systems theory, provides a precise definition of dynamical equivalence, offering a possible route to address this need. However, topological conjugacies have historically been challenging to compute. By leveraging advances in Koopman operator theory, we develop a framework for identifying conjugate and non-conjugate training dynamics. To validate our approach, we demonstrate that comparing Koopman eigenvalues can correctly identify a known equivalence between online mirror descent and online gradient descent. We then utilize our approach to: (a) identify non-conjugate training dynamics between shallow and wide fully connected neural networks; (b) characterize the early phase of training dynamics in convolutional neural networks; (c) uncover non-conjugate training dynamics in Transformers that do and do not undergo grokking. Our results, across a range of DNN architectures, illustrate the flexibility of our framework and highlight its potential for shedding new light on training dynamics.
@@ 1 Introduction
The analysis and experimentation of deep neural network (DNN) training continues to uncover new – and in some cases, surprising – phenomena. By changing the architecture, optimization hyper-parameters, and/or initialization, it is possible to identify regimes in which DNN parameters evolve along trajectories (in parameter space) with linear dynamics [1, 2], low-dimensional dynamics [3], correlated dynamics [4], lazy/rich dynamics [5, 6], and oscillatory dynamics [7, 8]. In some cases, the training dynamics have been linked with the performance of the trained model [7, 9, 10], providing new insight in DNN generalization. Additionally, detailed understanding of the dynamics has led to improvements in training efficiency [4, 11], demonstrating the practical implications such work can provide.
To obtain a more complete picture of DNN training, it is necessary to have a method by which equivalent dynamics can be identified and distinguished from other, non-equivalent dynamics. The construction of equivalence classes, which has fundamentally shaped the study of complex systems in other domains (e.g., phase transitions [12], bifurcations [13], defects in materials [14]), would advance the understanding of how architecture, optimization hyper-parameters, and initialization shape DNN training and could be leveraged to search for new phenomena. However, identifying equivalent and non-equivalent training dynamics is challenged by the need for methods that:
  • Go beyond the coarseness of loss. While useful as metrics, training and test loss can be shaped by non-dynamical features of the training (e.g., different initializations, different number of hidden units). Thus, different losses is neither necessary nor sufficient to conclude non-equivalent training dynamics.
  • Respect permutation symmetry. DNNs are invariant to the re-ordering, within layers, of their hidden units [15]. Thus, identifying that DNN parameters evolve along trajectories that occupy distinct parts of parameter space is not sufficient to conclude non-equivalent dynamics [16, 17].
We propose to use topological conjugacy [18], a notion of dynamical equivalence developed in the field of dynamical systems theory (Sec. 3.1), to address these limitations. Historically, topological conjugacy has been difficult to compute [19], especially when the equations governing the dynamical systems under study are not known. However, recent advances in Koopman operator theory [20, 21, 22] (Sec. 3.2) have enabled the identification of topological conjugacies from data [23] (Sec. 3.3). We explore the potential of this Koopman-based approach for identifying topological conjugacies in the domain of DNN training, finding that it is able to:
  • Recover a known nonlinear topological conjugacy between the training dynamics of online mirror descent and online gradient descent [24, 25, 26] (Sec. 4.1);
@@
  • Demonstrate the existence of conjugate training dynamics across different random initializations of FCNs [16] (Appendix C.4);
@@
  • Uncover non-conjugate training dynamics across Transformers that do, and that do not undergo delayed generalization (i.e., “grokking”) [28, 29] (Sec. 4.4).
That the same framework can be used across a number of DNN architectures to study a variety of dynamical phenomena during training demonstrates the generality of the approach. We conclude by discussing how it can be further improved to enable greater resolution of equivalent dynamics, and how it can be used to shed greater light on DNN training (Sec. 5).
@@ 2.1 Identification of DNN training dynamics phenomena
Analytical results have been obtained for the DNN training dynamics of shallow student-teacher [30, 31] and infinite width [1, 2] networks. For modern architectures (e.g. CNNs, Transformers), the training dynamics have been probed via analysis of computational experiments. Application of dimensionality reduction has led to the observation that parameters are quickly constrained to being optimized along low-dimensional subspaces of the high-dimensional parameter space [3, 4]. Inspection of losses, parameter and gradient magnitudes, etc. led to the identification of several transitions in the training dynamics of CNNs during the initial few epochs [27]. While insightful, this prior work cannot – except at a coarse-grained level – be used to determine whether the dynamics associated with training different DNN models (or training the same DNN model with different choices in hyper-parameters or initialization) are equivalent.
@@ 2.2 Koopman operator theory applied to DNN training
Data-driven implementations of Koopman operator theory have been used to model the dynamics of DNN training [32, 33, 34]. Because of the linearity of the learned Koopman models (Sec. 3.2), using them in place of standard gradient-based methods has led to reductions in computational costs associated with DNN training. Koopman-based methods have additionally been used to meta-learn optimizers for DNNs [35, 36]. The ability of Koopman models to capture features of training dynamics has been leveraged to develop new methods for pruning DNN parameters [37, 38] and new adaptive training methods [39]. While this prior work has demonstrated that accurate Koopman operator representations of the nonlinear training dynamics can be extracted, none have utilized the theory to identify topological conjugacies.
@@ 3.1 Topological conjugacy
Given two discrete-time dynamical maps [^1] For the sake of space, we describe only discrete-time dynamical systems, but the theory extends to continuous time dynamical systems. $T_{1}:X\rightarrow X$ and $T_{2}:Y\rightarrow Y$, a natural question to ask is whether they induce equivalent dynamical behavior. There are various possibilities for defining equivalence, but dynamical systems theory has made use of the notion of topological conjugacy [18] to identify when a smooth invertible mapping can be used to transform trajectories of $T_{1}$ to those of $T_{2}$ (and vice versa). Formally, $T_{1}$ and $T_{2}$ are said to be topologically conjugate if there exists a homeomorphism, $h:X\rightarrow Y$, such that
@@
It is straightforward to identify and construct conjugacies for linear systems. Let $X=Y=\mathbb{R}^{n}$, and let $T_{1}=A$ and $T_{2}=B$, where $A,B\in\mathbb{R}^{n\times n}$. These describe linear dynamical systems, as $x(t+1)=Ax(t)$ and $y(t+1)=By(t)$. In this setting, $A$ and $B$ are conjugate if there exists an $H\in\mathbb{R}^{n\times n}$, such that $y(t)=Hx(t)$ and $A=H^{-1}BH$. This can happen if and only if the eigenvalues of $A$ are the same as the eigenvalues of $B$. Thus, for linear systems, topological conjugacy can be used to construct equivalence classes, partitioning the space of dynamical systems into families of matrices that have the same spectra. However, for nonlinear systems, it is challenging to prove the existence or non-existence of conjugacies [19], limiting its use as a tool. In addition, historically it has not been possible to compute topological conjugacies for systems where the underlying dynamics are not analytically known.
@@ 3.2 Koopman mode decomposition
Figure 1: Schematic of Koopman operator theory-based identification of conjugate dynamical systems. (A) By lifting nonlinear dynamics from a finite dimensional state-space to an infinite dimensional function space, a linear representation can be achieved (from which a finite dimensional approximation can be obtained). (B) The linearity of the Koopman operator enables a mode decomposition, which includes Koopman eigenvalues (orange), eigenfunctions (green), and modes (blue). (C) Dynamical systems with the same Koopman eigenvalues are topologically conjugate.
@@
In cases of chaotic dynamics, a representation by a finite number of Koopman modes is not achievable. Such systems are said to have continuous spectra. In order for a DNN training algorithm to be useful, it likely must avoid chaotic behavior. Therefore, we focus on training dynamics where Eq. 4 is assumed to be valid.
@@
An important feature of the Koopman eigenvalues is that they are invariant to permutations of the ordering of state-space variables. Let $x=[x_{1},...,x_{n}]$ and $\tilde{x}=[x_{\sigma(1)},...,x_{\sigma(n)}]$, where
$\sigma:\{1,...,n\}\rightarrow\{1,...,n\}$ is a permutation and $\rho_{\sigma}:x\rightarrow\tilde{x}$ is the permutation mapping. That is, $\tilde{x}$ is equivalent to $x$ via a re-ordering of its labels. In this case, the action of the Koopman operator is
$U^{t}\tilde{g}(\tilde{x})=\sum_{i=1}^{N}\lambda_{i}^{t}\tilde{\phi}_{i}(\tilde{x})v_{i},$ (5)
where $\tilde{g}(\tilde{x})=g[\rho_{\sigma}^{-1}(\tilde{x})]$ and $\tilde{\phi}(\tilde{x})=\phi[\rho_{\sigma}^{-1}(\tilde{x})]$. Thus, the Koopman eigenvalues are the same as they were for the non-permuted system. We note that the Koopman spectrum is the same for other invariances that are known to exist in DNNs, such as rescaling (of cascaded linear layers) and rotations (of query and key projections used in attention in Transformers) [41]. This makes it a generally powerful approach for studying DNN training dynamics.
While Eq. 4 is true for deterministic dynamical systems and does not hold for training via stochastic gradient descent (SGD), we believe it is still to appropriate to compute the KMD from weight trajectories for two reasons. First, theoretical work has expanded the notion of Koopman operator theory to stochastic dynamical systems [42] and defined Eq. 4 in terms of the expectation of the dynamics. This suggests that the KMD associated with SGD training will be able to inform us of the “average” dynamics during training. This will be useful to comparing different network behaviors. And second, prior work computing KMD on DNN training has found it able to sufficiently approximate the training dynamics so as to allow for the Koopman operator to be used to optimize [32, 34] and sparsify [38] DNNs. This suggests KMD can capture important aspects of the training.
Many numerical methods have been developed for approximating the KMD from data. This has enabled its successful application as a tool for spatio-temporal decomposition in providing insight into complex, real-world dynamical systems [43, 44, 45, 46, 47]. Dynamic mode decomposition (DMD) [48, 49], the most popular of these methods, has spawned many variants [50, 51, 52, 53, 54]. In general, DMD-based approaches collect $T+1$ snapshots of data $x\in\mathbb{R}^{n}$, construct data matrices $Z=[x(0),...,x(T-1)]$ and $Z^{\prime}=[x(1),...,x(T)]$, where $Z,Z^{\prime}\in\mathbb{R}^{n\times(T+1)}$, and then approximate the Koopman operator by
@@
where $\dagger$ denotes the pseudo-inverse. Utilizing dictionaries with nonlinear functions [50] has led to improved results, demonstrating how usage of the underlying Koopman operator theory can enhance the capture of complex dynamics. In addition, leveraging Takens’ Embedding Theorem [55] and using time-delayed observables has proved to be a generally powerful approach for approximating Koopman eigenvalues [44, 45, 51], an approach we make use of (Sec. 4).
@@ 3.3 Equivalent Koopman spectra implies topological conjugacy
Given that KMD provides a linear representation of nonlinear dynamical systems, identifying topological conjugacies through matching eigenvalues (Sec. 3.1) again becomes viable. Indeed, it has been proven that two discrete-time dynamical maps $T_{1}$ and $T_{2}$, each in the basin of attraction of a stable fixed point, are topologically conjugate if and only if the Koopman eigenvalues of the associated Koopman operators, $U_{1}$ and $U_{2}$, are the same [23] (Fig. 1C). That is, a topological conjugacy exists if and only if
@@
where $\lambda^{(1)}$ and $\lambda^{(2)}$ correspond to the eigenvalues associated with $U_{1}$ and $U_{2}$, respectively, and $N$ is the number of Koopman modes. When the dynamical systems under study have continuous spectra, Eq. 7 does not imply topological conjugacy. As noted earlier, we do not believe this to be a major limitation when studying meaningful training dynamics. However, recent work has suggested that topological conjugacies may still be identifiable in the case of continuous spectra by using extensions of Koopman operator theory [56]. We believe this will be a fruitful direction for future work.
When the number of Koopman eigenvalues of $U_{1}$ is larger than the number of Koopman eigenvalues of $U_{2}$, the strict equivalence of Eq. 7 cannot be satisfied. However, there may exist a smooth, but non-invertible mapping $h$ from $X$ onto $Y$. In such a case, $T_{1}$ and $T_{2}$ are said to be semi-conjugate, and this can be identified when $\{\lambda_{i}^{(2)}\}_{i=1}^{N_{2}}\subset\{\lambda_{j}^{(1)}\}_{j=1}^{N_{1}}$, where $N_{2}<N_{1}$ are the number of Koopman eigenvalues of $U_{1}$ and $U_{2}$, respectively.
Computing the KMD from data is unlikely to yield the same exact Koopman eigenvalues for conjugate dynamical systems, due to the presence of noise and finite sampling. Therefore, a method for computing the distance between eigenvalues is necessary when making comparisons. Here, we make use of the Wasserstein distance [57, 58], a metric developed in the context of optimal transport that quantifies how much one distribution must be changed to match another. This notion of “distance” is important as the Koopman eigenvalues correspond to time-scales and we expect dynamical systems with increasingly large differences between their eigenvalues will have increasingly large differences in their dynamical behavior [^2] Note that other metrics, such as the KL-divergence, may not capture this distinction.. In the case where a small, finite number of Koopman eigenvalues are computed (which can be achieved, even for systems with a large number of observables, by performing dimensionality reduction or residual based pruning of modes [53]), the Wasserstein distance can be efficiently computed by using linear sum assignment.
@@ 4.1 Identifying conjugate optimizers
Figure 2: Conjugacy between online mirror descent and online gradient descent is identifiable from Koopman spectra. (A) Comparing example trajectories of variables optimized via OMD ($x_{1},x_{2}$), OGD ($u_{1},u_{2}$), and BM ($z_{1},z_{2}$), the existence of a conjugacy between OMD and OGD is not obvious. (B) Similarly, the existence of a conjugacy is not apparent when looking at the loss incurred by using OMD and OGD. (C) Comparing the Koopman eigenvalues associated with optimizing using OMD, OGD, and BM correctly identifies the existence of a conjugacy between OMD and OGD, and the lack of a conjugacy between OMD/OGD and BM. The function optimized is in all subfigures is $f(x)=\sum\tan(x)$.
We compute the KMD associated with optimization using OMD and OGD by considering trajectories of both, from many initial conditions, and compare the resulting Koopman eigenvalues (Appendix A.4). We find high overlap between the two spectra (Fig. 2C, red and black dots). Additionally, the two sets of eigenvalues have the same structure. Namely, they consist only of real, positive eigenvalues. In contrast, the bisection method (BM), another optimization algorithm that is not conjugate to OMD or OGD (Appendix A.2), has associated spectra that are complex (Fig. 2C, light blue dots). Performing a randomized shuffle of the eigenvalues between algorithms (Appendix B), we find that $25\%$ of the shuffles between OMD and OGD eigenvalues result in Wasserstein distance greater than or equal the true Wasserstein distance. This suggests the distributions are not statistically significantly distinct. However, $0\%$ of the shuffles have Wasserstein distance greater than or equal to the true Wasserstein distance for OMD and BM, and OGD and BM, respectively. This provides evidence that the spectra of OMD/OGD and BM are statistically significantly distinct.
Similar results are obtained when applying KMD to OMD and OGD optimization of a different function (Fig. S1). Collectively, these results demonstrate that the Koopman-based spectral identification of topological conjugacies can successfully recover a known equivalence and provide support that it can be used more broadly in uncovering equivalences in DNN training dynamics.
4.2 Identifying the effect of width on fully connected neural network training
To start exploring the potential of our framework for identifying topological conjugacies in DNN training, we begin with a small-scale example. Namely, we consider a fully connected neural network (FCN) with only a single hidden layer, trained on MNIST (Appendix C.1). Consistent with other architectures, we find that the wider the FCN (i.e., the more units in the hidden layer), the better the performance and the lower the loss (Fig. 3A). Whether this is due to an increase in capacity, with more hidden units enabling a more refined solution, or whether this is due to a change in the training dynamics, leading to a better traversal of the loss landscape, is – at this point – unclear.